Journalist Report – February 6th

Crew 274 Journalist Report 2-6-2023

Sol: 8

Author: Tony DiBernardo, Crew Journalist

Title: Media Day

Crew 274 woke up this morning at 7:00am to the song “Africa” by Toto. The crew got their surveys done early because they had a big day ahead of them. Unlike our past Journalist Reports, this report will not include sections for EVA, Hab Design, of Engineering, Green Hab, or Stress Mitigation.

Today, the crew broke simulation for a media and outreach day. We conducted eleven broadcasts to Family members, the APUS Graduate Studies Office, 3 classrooms from Waugh Chapel Elementary School, a class from Enemy Swim Day School from Waubay, South Dakota, and teams from SpaceX and the Vance Brand Squadron of the Civil Air Patrol.

In total, we spoke with over 170 adults and elementary school students. They asked incredibly insightful questions like:

  • What is the temperature on Mars?

  • How long does it take to get there?

  • What happens when you enter a black hole?

  • How do you “go potty” in space?

  • What do you eat in space?

  • Could you get along with each other on a real mission to Mars?

  • Are you having fun?

  • Why do you go to Mars and not a different planet?

  • How long does it take for a plant to grow in space?

  • Have you ever seen an alien or alien spaceship?

The crew also taught them about the whole analog experience, visual star processing, solar weather processing, how space weather impacts orbital assets, and how it would be helpful for Martians on site to have real-time access to solar weather. The team had a blast communicating their science to family, friends and the general public, but we look forward to getting back into our routine, our research, and finishing the last half of our mission strong.

GreenHab Report – February 6th



GreenHab Officer: Tyler Hines

Environmental control: heater

Average temperatures: 75.05 F

Hours of supplemental light: 4 hours

Daily water usage for crops: 12.68 gallons

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank: 87.37 gallons

Time(s) of watering for crops: 11:23 am

Changes to crops: N/A

Narrative: Continued to conduct watering of plants and crops during daily GreenHab operations and finalized the list of another potential harvest while also cleaning and trimming any notable debris and or dead leaves from any crops and plants.

Harvest: N/A

Support/supplies needed: N/A

Astronomy Report – February 6th

Astronomy Report

Name: Salina Peña and Noah Loy

Crew: 274

Date: 02/06/2023


Robotic Telescope Requested (Choose one) MDRS-14 MDRS-WF

Objects Viewed: None

Problems Encountered: None


Solar Features Observed: None

Problems Encountered: None due to cloud interference.

Sol Summary Report – February 6th

Crew 274 Sol Summary Report 02-06-2023

Sol: 8

Summary Title: Public Outreach Day

Author’s name: Sarah E. Guthrie (“Ceres”), Commander

Mission Status: Nominal

Sol Activity Summary: The crew hosted a public outreach day with multiple briefings that filled the day. The crew completed sessions with elementary and middle schools from the Dakota Nation and Maryland, SpaceX, Civil Air Patrol, university leadership, families, friends, and colleagues. The crew completed surveys and prepared for possible EVA tomorrow.

Look Ahead Plan: The crew looks forward to the emergency rescue of KURT and the assessing the cache system (pending approval).

Anomalies in work: None

Weather: Sunny, 1C High, -6C Low; partly sunny, windy (16 mph)evening clouds

Crew Physical Status: Excited

EVA: EVA #6 completed for logistics cache assessment – successful

Reports to be filed: Sol Summary, Operations, EVA Request, GHO Report, and Astronomy Report.

Support Requested: None

Operations Report – 6th

Crew 274 Operations Report 2/6/2023

SOL: 8

Name of person filing report: Alexis Lojek

Non-nominal systems: Toilet

Notes on non-nominal systems: Mission Support will install a new toilet holding tank between crews.


Spirit rover used: No

Hours: 208.0

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

Opportunity rover used: No

Hours: 112.2

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

Curiosity rover used: No

Hours: 220.3

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

Perseverance rover used: Yes

Hours: 255.2

Beginning charge: 100%

Ending charge: N/A

Currently charging: No

General notes on rovers: No rovers were utilized today.

Summary of Hab operations: Toilet was drained into the septic system this morning.

WATER USE: 39 gallons as of 4:09 pm.

Water (static tank): 249 gallons remaining

Static tank pipe heater (On or off): On

Static tank heater (On or off): On

Toilet tank emptied: Yes

Summary of internet: Nominal

Summary of suits and radios: Suits are in nominal condition and charging. Radios were charged yesterday and are off chargers. All conditions normal.

Summary of GreenHab operations: GHO Hines watered plants from 11:23 – 12:00 pm.

WATER USE: 12.68 gallons used

Heater: On

Supplemental light: 4 hours (10pm-2am)

Harvest: No

Summary of ScienceDome operations: GHO Officer Tyler Hines’ experiment is continuing to be monitored; Cress microgreens are almost completely sprouted.

Dual split: Still on a timer from 10pm to 7am and under Mission Support supervision.

Summary of RAM operations: No operations were conducted in the RAM.

Summary of any observatory issues: No observatory issues.

Summary of health and safety issues: N/A

Questions, concerns, and requests to Mission Support: N/A

Copyright © The Mars Society. All rights reserved. | Main Site